OpenStudy (anonymous):

multiplying rational functions and equations: p^6*q^2/7r^3 * -3p^2/r

OpenStudy (saifoo.khan):

\[\frac{p^6 q^2 }{7r^3} \times -\frac{3p^2}{r}\] \[\large -\frac{3p^8 q^2}{7r^4}\]

OpenStudy (mertsj):

\[\frac{p^6q^2}{7r^3}\times\frac{-3p^2}{r}=\frac{-3p^8q^2}{7r^4}\]